A stock just paid a dividend of $2.59. The dividend is expected to grow at 20.02% for five years and then grow at 4.69% thereafter. The required return on the stock is 11.99%. What is the value of the stock?
Answer #1

Year 1 dividend = 2.59 (1 + 20.02%) = 3.108518

Year 2 dividend = 3.108518 (1 + 20.02%) = 3.730843

Year 3 dividend = 3.730843 (1 + 20.02%) = 4.477758

Year 4 dividend = 4.477758 (1 + 20.02%) = 5.374205

Year 5 dividend = 5.374205 (1 + 20.02%) = 6.450121

Year 6 dividend = 6.450121 (1 + 4.69%) = 6.752632

Value at year 5 = D6 / required rate - growth rate

Value at year 5 = 6.752632 / 0.1199 - 0.0469

Value at year 5 = 6.752632 / 0.073

Value at year 5 = 92.501808

Value of stock = 3.108518 / (1 + 0.1199)1 + 3.730843 / (1 + 0.1199)2 + 4.477758 / (1 + 0.1199)3 + 5.374205 / (1 + 0.1199)4 + 6.450121 / (1 + 0.1199)5 + 92.501808 / (1 + 0.1199)5

Value of stock = $68.53